# Diffusion Suction Grasping with Large-Scale Parcel Dataset

### v1

@@ III-C Dataset Details
To the best of our knowledge, this comprehensive dataset is the first large-scale synthetic suction grasping dataset in parcel stacked scenes, called Parcel-Suction-Dataset. Specifically, it contains 500 cycles and each cycle contains 50 scenes. There are a total of 25 thousand scenes included. To cover the entire scene, 16384 extracted annotations are sampled from each scene using farthest point sampling. Each scene includes RGB images, segmentation images, depth maps, scene point clouds, 6D object poses, 6D object bounding boxes and camera matrices. In addition, the dataset is divided into a training set and a testing set. The training set contains 480 cycles, and the test set contains 20 cycles. The training set contains 100 objects, and the test set contains 13 other objects that does not appear in the training set. Notably, this Self-Parcel-Suction-Labeling framework is intended to serve as a synthetic benchmark for parcel suction grasping.
@@ IV Diffusion-Suction Model
Figure 3: Overview of Diffusion-Suction architecture. Based on the diffusion model, Diffusion-Suction learns an iterative denoising process to denoise random noise into suction grasping score map with the guidance of input point clouds. It uses PointNet++ to extract features as condition guidance. It consists of a novel and light-weighted Pointcloud Conditioned Denoising Block. During the merge stage, we apply Non-Maximal Suppression as a post-processing step. The red spheres represent the predicted score ($S>0.8$). It can be observed that the reverse process is a gradual shift from random noise to suction grasping score.
@@ IV-A Task Reformulation
Preliminaries. Diffusion models [16, 23] are a class of generative models parametrized by a Markov chain, which can gradually transform a simple distribution to a structured complex distribution. These models involve two processes. The first is diffusion process $q$, which can be defined as:
@@
Denoising as Suction Grasping. Given a stacked scene where parcels are stacked randomly into a pile, we are interested in detecting the suction grasping pose. The pose is defined as:
@@
where $t$ represents suction cup position on the object surface, and $n$ represents the orientation of the suction cup. We define suction grasping prediction task as, given the point clouds of the parcel stacked scenes $P$, the algorithm predicts a feasible suction grasping poses $V$. The orientation of the suction grasping $n$ usually follows the direction of the normal vector on the object surface, and the normal vector can be calculated directly. Therefore, the suction grasping task is actually transformed into finding a feasible suction grasping pose $V$ with a high confidence score $x$ in the scene. Subsequently, the diffusion process will occur in the confidence score space. In this paper, we reformulate the suction grasping prediction as a 3D visual-condition guided denoising process, which can be defined as:
@@
where the reverse process model ${\mu_{\theta}}({x_{t}},t,P)$ is trained to refine suction grasping score from latent ${x_{t}}$ to ${x_{t-1}}$ conditioned on the scene point cloud ${P}$.
@@ IV-B Architecture
The diffusion model generates data samples progressively, requiring the model ${\mu_{\theta}}({x_{t}},t,P)$ runs multiple times during the inference stage. However, directly applying the model ${\mu_{\theta}}({x_{t}},t,P)$ on the raw scene point clouds at each iterative step significantly increases the computational overhead. To resolve this problem, we propose to separate the entire model into two parts: point clouds encoder and suction grasping decoder, as shown in Fig. 3. The point clouds encoder runs only once to extract the deep feature representation from the input scene point clouds ${P}$. Then the suction grasping decoder takes this deep feature as visual-condition guided, instead of the raw point clouds ${P}$, to progressively refine the suction grasping prediction score from the noisy ${x_{t}}$. With this simple and efficient structure, DiffusionSuction reduces the inference computational overhead.
Point cloud encoder. It takes the original point clouds of the parcel stacked scenes as input and applies a feedforward network for feature extraction. For this purpose, we employ PointNet++[24] serving as the backbone and there are other alternative backbones, e.g. DGCNN [25], Pointnet [26] and MinkowskiEngine [27]. The extracted point-wise guidance $F_{p}$ has a size of $N_{p}\times N_{f}$. Calculating object surface normals can effectively extract the main 3D structure information of the object and provide effective prior information without the training process. Therefore, we use Open3D [28] to directly estimate normals and then fuse 3D normals prior feature with the point clouds to obtain accurate prediction.
Suction grasping decoder. Neural network model ${\mu_{\theta}}({x_{t}},t,P)$ takes available 3D visual information feature from parcel stacked scene and iteratively refines suction grasping noisy map ${x_{t}}$. We introduce light-weighted Pointcloud Conditioned Denoising Block (PCDB) to achieve this process, which efficiently reuses the shared parameters during the multi-step reverse diffusion process. Specifically, as shown in Fig. 3, suction grasping noisy map ${x_{t}}$, time steps embedding $t$ and visual-condition guided $F_{p}$ are sent to different MPLs for powerful feature representation. Afterwards, the three feature maps are combined via element-wise addition. To emphasize the significant features, we conduct attention on the channel and spatial information with the convolutional block attention module (CBAM) [29]. Inspired by the success of residual module [30], we conduct residual connection between the fused attention feature maps ${x_{t}}$ to obtain ${x_{t-1}}$. Our PCDB balances effectiveness and efficiency, refining suction grasping with few parameters.
@@ IV-C Training and Inference
We add Gaussian noise to the ground-truth suction grasping score,according to Equation (5). The noise scale is controlled by ${a_{t}}$, which adopts the monotonically cosine schedule for ${a_{t}}$ in different time steps. Trainable parameters are mainly the PCDB and point cloud feature extractors. The model is trained by minimizing the loss between diffusion suction grasping prediction and ground-truth, which is provided in Algorithm 1. The iteration number $1\leq t\leq T$ is sampled from a uniform distribution, and the epsilon $\boldsymbol{\epsilon}$ is sampled from a standard distribution. Notably, the scaling factor scale controls the signal-to-noise ratio (SNR) [31], which has a significant effect on the performance of the model. Therefore, we normalize and scale the range of ground-truth.
@@
The inference procedure of Diffusion-Suction is a multi-step denoising process from a Gaussian distribution. Given a raw scene point clouds as the condition input, the model progressively refines its predictions and starts with a noise map, as shown in Algorithm 2. Moreover, we apply the DDIM [11] inversion process for the improving inference process. Fig. 4 illustrates the denoising reverse process with 20 inference steps of Diffusion-Suction.
@@
Figure 4: The figure shows the denoising reverse process with 20 inference steps of Diffusion-Suction. The blue cylinders represents a suction pose and the color intensity indicates the confidence of suction pose. It can be observed that the reverse process is a gradual shift from pure noise to refined suction score.
@@ V Experiments
TABLE I: Quantitative evaluation on SuctionNet-1Billion. The evaluation results include Kinect and Realsense collection source.
@@
TABLE II: Quantitative evaluation of Suction Grasping on the Parcel-Suction-Dataset.
@@ V-A Datasets and evaluation metrics
To comprehensively evaluate the suction grasping performance of our method in stacked parcel scenes, we select the Parcel-Suction-Dataset. There are a total of 25 thousand annotated scenes included. Moreover, we also choose to conduct testing on the public datasets SuctionNet-1Billion [3] to demonstrate the effectiveness of our framework. This dataset contains 190 scenes and we collect 256 perspectives for each scene. The dataset is divided into three subsets to test performance in different scenes: seen, similar, and novel.
We evaluate the method’s performance by calculating the Average Precision metric. Specifically, we strictly follow the evaluation protocol in proposed method [3] and calculate the seal score and wrench score for each predicted result online. If the product of the seal score and the wrench score exceeds the threshold of 0.4 or 0.8, the predicted suction pose is considered correct. To evaluate the model’s performance in real-world cluttered scenes, we also select the prediction grasping pose with the highest confidence for evaluation. When evaluating the Parcel-Suction-Dataset, we calculate visibility scores for online evaluation.
@@ V-B Evaluation on Parcel-Suction-Dataset
To validate the performance of our proposed approach on the Parcel-Suction-Dataset, we compare Diffusion-Suction against several state-of-the-art approaches, as shown in Table II. As we can observe, our proposed approach achieves state-of-the-art results. It outperforms other methods wtih an improvement of +32% on average precision the in the top-50 metric. Moreover, our Diffusion-Suction achieves great improvement in the Top-1 metric, indicating its ability to generate the best suction grasping candidate in cluttered scenes, which is crucial for real-world suction grasping applications. Notably, in terms of the ${AP_{0.8}}$ metric, our method significantly outperforms other approaches, showing the most substantial improvement. This indicates that most of suction grasping poses predicted by Diffusion-Suction have high scores. Fig. 5 illustrates the qualitative results of Diffusion-Suction on the cluttered and complex parcel scenes, showing the Top-50 suction grasps.
@@ V-C Evaluation on SuctionNet-1Billion
Table I summarizes the comparison results between our proposed approach and current suction grasping methods on public benchmark. Overall, our proposed approach outperforms three state-of-the-art approaches by a significant margin across both RealSense and Kinect collection source. Specifically, it’s important to note that our proposed approach demonstrates consistently better performance in all the top-50 benchmark compared wtih other methods. A crucial contributing factor is that our diffusion head with advanced 3D visual-condition feature extraction can better capture global information, leading to highly accurate prediction overall performance. In addition, our proposed approach outperforms other models in all the seen scenes and exhibits strong learning capability for seen objects. It shows that utilizing diffusion model on this task can achieve robustness and efficiency prediction in terms of the seen scenes. Our proposed approach also demonstrates competitive performance on the similar and novel scenes.
@@ V-D Ablation study
We present extensive ablation study to further reveal the properties of Diffusion-Suction and the ablation is conducted on the the Parcel-Suction-Dataset.
@@
Signal scale. We study the effect of different scaling factor, which controls the signal-to-noise ratio of the diffusion process, as shown in Table IV. Diffusion-Suction with the signal scaling factor of 0.5 demonstrates optimal performance. We explain that this is because Diffusion-Suction with a larger scaling factor results in more easier cases being preserved with the same time step.
Denoising step. Training and inference step not only affect inference speed, but also require high GPU memory consumption. To explore the balance between accuracy and efficiency, we change the training and inference step within the range of 5 to 20, as shown in Table V. Changing the inference step directly can result in a significant decrease in performance. However, keeping the training and inference step the same and reducing the number of step slightly decrease the performance, but it can accelerate the model speed.
@@
TABLE IV: Ablation study on different scale settings.
  Scale ${AP}$ ${AP_{0.8}}$ ${AP_{0.4}}$
Diffusion-Suction(0.001) 93.29 80.48 97.62
Diffusion-Suction(0.01) 93.00 79.50 97.59
Diffusion-Suction(0.1) 92.88 79.28 97.52
Diffusion-Suction(0.5) 94.42 85.04 97.65
Diffusion-Suction(1) 94.02 84.69 97.20
TABLE V: Ablation study on different training and inference step.

### v2

@@ III-C Dataset Details
To the best of our knowledge, this comprehensive dataset is the first large-scale synthetic suction grasping dataset in parcel stacked scenes, called Parcel-Suction-Dataset. Specifically, it contains 500 cycles and each cycle contains 50 scenes. There are a total of 25 thousand scenes included. To cover the entire scene, 16384 extracted annotations are sampled from each scene using farthest point sampling. Each scene includes RGB images, segmentation images, depth maps, scene point clouds, 6D object poses, 6D object bounding boxes and camera matrices. The training set contains 100 objects, and the test set contains 13 other objects that does not appear in the training set. Notably, this Self-Parcel-Suction-Labeling framework is intended to serve as a synthetic benchmark for parcel suction grasping.
@@ IV Diffusion-Suction Model
Figure 3: Overview of the Diffusion-Suction architecture. Diffusion-Suction learns an iterative denoising process to denoise random noise into suction grasping score map with the guidance of input point clouds. It uses PointNet++ to extract features as condition guidance. It consists of a novel and lightweight Pointcloud Conditioned Denoising Block. During the filtering stage, we apply Non-Maximal Suppression as a post-processing step. The red spheres represent the predicted scores ($S>0.8$). It can be observed that the reverse process is a gradual shift from random noise to suction grasping scores.
@@ IV-A Task Reformulation
Preliminaries. Diffusion models [17, 21] are a class of generative models parametrized by a Markov chain, which can gradually transform a simple distribution to a structured complex distribution. These models involve two processes. The first is the diffusion process $q$, which can be defined as:
@@
Denoising as Suction Grasping. Given a stacked scene where parcels are stacked randomly into a pile, we are interested in detecting the suction grasping poses. The poses are defined as:
@@
where $t$ represents the suction cup position on the object surface, and $n$ represents the orientation of the suction cup. We define suction grasping prediction task as follows: given the point clouds of the parcel-stacked scenes $P$, the algorithm predicts a set of feasible suction grasping poses $V$. The orientation of the suction grasping $n$ usually follows the direction of the normal vector on the object surface, and the normal vector can be calculated directly. Therefore, the suction grasping task is actually transformed into finding a set of feasible suction grasping poses $V$ with a high confidence score $x$ in the scene. Subsequently, the diffusion process will occur in the confidence score space. In this paper, we reformulate the suction grasping prediction as a 3D visual-condition guided denoising process, which can be defined as:
@@
where the reverse process model ${\mu_{\theta}}({x_{t}},t,P)$ is trained to refine suction grasping scores from latent ${x_{t}}$ to ${x_{t-1}}$ conditioned on the scene point cloud ${P}$.
@@ IV-B Architecture
The diffusion model generates data samples progressively, requiring the model ${\mu_{\theta}}({x_{t}},t,P)$ to run multiple times during the inference stage. However, directly applying the model ${\mu_{\theta}}({x_{t}},t,P)$ on the raw scene point clouds at each iterative step significantly increases the computational overhead. To resolve this problem, we propose to separate the entire model into two parts: point cloud encoder and suction grasping decoder, as shown in Fig. 3. The point clouds encoder runs only once to extract the deep feature representations from the input scene point clouds ${P}$. Then the suction grasping decoder takes these deep features as visual-condition guidance, instead of the raw point clouds ${P}$, to progressively refine the suction grasping prediction score from the noisy ${x_{t}}$. With this simple and efficient structure, Diffusion-Suction reduces the computational overhead during inference.
Point cloud encoder. It takes the original point clouds of the parcel stacked scenes as input and applies a feedforward network for feature extraction. For this purpose, we employ PointNet++[22] serving as the backbone and there are other alternative backbones, e.g. DGCNN [23], Pointnet [24] and MinkowskiEngine [25]. The extracted point-wise guidance $F_{p}$ has a size of $N_{p}\times N_{f}$. Calculating object surface normals can effectively extract the main 3D structure information of the object and provide effective prior information without a training process. Therefore, we use Open3D [26] to directly estimate normals and then fuse the 3D normals prior feature with the point clouds to obtain accurate predictions.
Suction grasping decoder. The neural network model ${\mu_{\theta}}({x_{t}},t,P)$ takes available 3D visual information features from parcel stacked scene and iteratively refines suction grasping noisy map ${x_{t}}$. We introduce a lightweight Pointcloud Conditioned Denoising Block (PCDB) to achieve this process, which efficiently reuses shared parameters during the multi-step reverse diffusion process. Specifically, as shown in Fig. 3, suction grasping noisy map ${x_{t}}$, time steps embedding $t$ and visual-condition guided $F_{p}$ are sent to different MLPs for powerful feature representation. Afterwards, the three feature maps are combined via element-wise addition. To emphasize the significant features, we conduct attention on the channel and spatial information with the convolutional block attention module (CBAM) [27]. Inspired by the success of the residual module [28], we conduct residual connection between the fused attention feature maps ${x_{t}}$ to obtain ${x_{t-1}}$. Our PCDB balances effectiveness and efficiency, refining suction grasping with few parameters.
@@ IV-C Training and Inference
We add Gaussian noise to the ground-truth suction grasping score according to Equation (5). The noise scale is controlled by ${a_{t}}$, which adopts the monotonically cosine schedule for ${a_{t}}$ in different time steps. The trainable parameters primarily include the PCDB and point cloud feature extractors. The model is trained by minimizing the loss between the diffusion suction grasping prediction and ground-truth, which is provided in Algorithm 1. The iteration number $1\leq t\leq T$ is sampled from a uniform distribution, and the epsilon $\boldsymbol{\epsilon}$ is sampled from a standard distribution. Notably, the scaling factor scale controls the signal-to-noise ratio (SNR) [29], which has a significant effect on the performance of the model. Therefore, we normalize and scale the range of the ground-truth. We found that Diffusion-Suction with a 0.5 signal scaling factor achieves optimal performance because larger scaling factors preserve a higher number of easier cases within the same time step.
@@
The inference procedure of Diffusion-Suction is a multi-step denoising process from a Gaussian distribution. Given raw scene point clouds as the condition input, the model progressively refines its predictions, beginning with a noise map, as shown in Algorithm 2. Moreover, we apply the DDIM [10] inversion process to improve the inference process. Fig. 4 illustrates the denoising reverse process with 20 inference steps of Diffusion-Suction.
@@
Figure 4: The figure shows the denoising reverse process with 20 inference steps. A blue cylinder represents a suction pose and the color intensity indicates the confidence of suction pose. It can be observed that the reverse process is a gradual shift from pure noise to refined suction score.
@@ V Experiments
TABLE I: Quantitative evaluation on SuctionNet-1Billion. Evaluation results include Kinect/RealSense collection sources, where data before/after the slash correspond respectively.
@@
TABLE II: Quantitative evaluation on the Parcel-Suction-Dataset.
@@ V-A Datasets and evaluation metrics
We evaluate our method on two datasets: our proposed Parcel-Suction-Dataset containing 25 thousand annotated stacked parcel scenes, and the public SuctionNet-1Billion benchmark [3] to demonstrate the framework’s effectiveness. This dataset contains 190 scenes and we collect 256 perspectives for each scene. The dataset is divided into three subsets to test performance in different scenes: seen, similar, and novel.
For performance evaluation, we adopt the Average Precision (AP) metric following the protocol in [3], which computes seal and wrench scores for each prediction online. If the product of the seal score and the wrench score exceeds the threshold of 0.4 or 0.8, the predicted suction pose is considered correct. When evaluating the Parcel-Suction-Dataset, we calculate visibility scores for online evaluation.
@@ V-B Evaluation on Parcel-Suction-Dataset
To validate the performance of our proposed approach on the Parcel-Suction-Dataset, we compare Diffusion-Suction against several state-of-the-art approaches, as shown in Table II. As we can observe, our proposed approach achieves state-of-the-art results. It outperforms other methods with an improvement of +32% on average precision the in the top-50 metric. Moreover, our Diffusion-Suction achieves great improvement in the Top-1 metric, indicating its ability to generate the best suction grasping candidate in cluttered scenes, which is crucial for real-world suction grasping applications. Notably, in terms of the ${AP_{0.8}}$ metric, our method significantly outperforms other approaches, showing the most substantial improvement. This indicates that most of suction grasping poses predicted by Diffusion-Suction have high scores. Fig. 5 illustrates the qualitative results of Diffusion-Suction on the cluttered and complex parcel scenes, showing the Top-50 suction grasps.
@@ V-C Evaluation on SuctionNet-1Billion
Table I summarizes the comparison results between our proposed approach and current suction grasping methods on public benchmark. Overall, our proposed approach significantly outperforms three state-of-the-art approaches across both RealSense and Kinect data source. Specifically, it’s important to note that our proposed approach demonstrates consistently better performance in all the top-50 benchmark compared with other methods. This superior performance can be attributed to our diffusion head’s advanced 3D visual-condition feature extraction capability, which better captures global information and leads to more accurate predictions. In addition, our proposed approach outperforms other models in all the seen scenes and exhibits strong learning capability for seen objects. It shows that utilizing diffusion model on this task can achieve robustness and efficiency prediction in terms of the seen scenes. Our proposed approach also demonstrates competitive performance on the similar and novel scenes.
@@ V-D Ablation study
We conduct extensive ablation studies on the Parcel-Suction-Dataset to analyze the key components and properties of Diffusion-Suction.
@@
Denoising step. Training and inference step not only affect inference speed, but also require high GPU memory consumption. To explore the balance between accuracy and efficiency, we change the training and inference step, as shown in Table IV. Changing the inference step directly can result in a significant decrease in performance. However, keeping the training and inference step the same and reducing the number of step slightly decrease the performance, but it can accelerate the model speed.
@@
TABLE IV: Ablation study on different training and inference step.
@@
TABLE V: Robotic grasping experiments on the real-world parcel scenes with the proposed approach Diffusion-Suction.
Object Scattered Scene Stacked Scene
${R_{grasp}}$ ${R_{object}}$ ${R_{grasp}}$ ${R_{object}}$
Rectangular 96.43 100.0 93.75 100.0
Planar 100.0 100.0 94.64 100.0
Cylindrical 81.36 90.83 78.69 100.0
Mixed 94.44 97.62 93.33 100.0
V-E Real World Experiment
Real-robot experiments are conducted using the Kinova Gen3 robot and a Mech-Eye industrial 3D camera, as shown in Fig. 6. To rigorously evaluate robustness across diverse grasping conditions, we conducted tests in diverse scenarios involving four types of objects: rectangular, flat, cylindrical, and mixed (a combination of the former three). Additionally, we configured scenes according to difficulty levels as scattered and stacked Scenes. Two metrics are adopted: 1) grasping success rate ${R_{grasp}}$ = Successful grasps/Total attempts, 2) object clearance rate ${R_{object}}$ = Cleared objects /Total objects. The experiment are terminated if the robot failed to grasp any object in three consecutive attempts. Quantitative results, as shown in Table V, demonstrate our method’s effectiveness in real-world parcel scenes.
Figure 6: Real-world grasp experiment setup.